OpenStudy (anonymous):

Ramona deposited $4,190.51 into a savings account with an interest rate of 5.2% compounded twice a year. About how long will it take for the account to be worth $9,000? 33 years, 5 months 29 years, 9 months 14 years, 11 months 20 years, 8 months 9,000=4,190.51(1+0.052/2)^2t 4,809.49= 1.026^2t

OpenStudy (amistre64):

if we just let the exponent reference 6month time periods, we can ignore the 2 on it

OpenStudy (amistre64):

A = P(1+r/2)^t A/P = (1+r/2)^t ln(A/P) = t ln((1+r/2)) ln(A/P)/ln(1+r/2) = t and 1t is equal to 6 months, or half a year